# SQL-GEN: Bridging the Dialect Gap for Text-to-SQL Via Synthetic Data And Model Merging

### v1

@@ Abstract
Text-to-SQL systems, which convert natural language queries into SQL commands, have seen significant progress primarily for the SQLite dialect. However, adapting these systems to other SQL dialects like BigQuery and PostgreSQL remains a challenge due to the diversity in SQL syntax and functions. We introduce SQL-GEN, a framework for generating high-quality dialect-specific synthetic data guided by dialect-specific tutorials, and demonstrate its effectiveness in creating training datasets for multiple dialects. Our approach significantly improves performance, by up to 20%, over previous methods and reduces the gap with large-scale human-annotated datasets. Moreover, combining our synthetic data with human-annotated data provides additional performance boosts of 3.3% to 5.6%. We also introduce a novel Mixture of Experts (MoE) initialization method that integrates dialect-specific models into a unified system by merging self-attention layers and initializing the gates with dialect-specific keywords, further enhancing performance across different SQL dialects.
@@ 1 Introduction
Text-to-SQL systems translate natural language questions into executable SQL queries, enabling users to interact with databases using natural language. This transformation is crucial as it links the intuitive nature of human communication with the structured precision of SQL, the standard language for querying databases (Androutsopoulos et al., 1995; Hristidis et al., 2003; Li & Jagadish, 2014). Text-to-SQL plays a significant role for conversational agents, empowering them to process complex queries within large-scale databases efficiently (Yu et al., 2019; Gu et al., 2022; Pérez-Mercado et al., 2023). Such systems serve as a copilot for data science professionals to enhance productivity, beyond being valuable for non-technical users who wish to derive business insights without SQL expertise (Li et al., 2023; Sun et al., 2023a, b; Wang et al., 2019).
SQL has been adopted by each database product (e.g. PostgreSQL, MySQL, and SQLite) to suit their specific needs. Despite their common foundations, these SQL dialects differ significantly in their syntax, functions, and capabilities, which even make the automated translation of queries across dialects a complex task that often requires human intervention (Zmigrod et al., 2024; Ngom & Kraska, 2024). Figure 1 exemplifies a question that can be answered with different SQL keywords across different dialects with their own unique keywords that are distinct from one another. Additionally in Section A.8, we provide some of the dialect specific keywords for BigQuery, PostgreSQL, and SQLite, which are not supported across all of them. In the realm of Text-to-SQL, most benchmarks are based on the SQLite dialect, chosen for its simplicity and self-contained nature (Li et al., 2024b; Yu et al., 2018b; Zhong et al., 2017; Chang et al., 2023; Gan et al., 2021). However, queries tailored to one dialect often fail in another – for instance as reported in Pourreza & Rafiei (2023), 32% of the queries from the Spider development set (Yu et al., 2018b) exhibit syntax errors when executed on a PostgreSQL database. This dialect dependency poses a significant challenge, as models trained on SQLite-specific syntax are prone to generating erroneous queries in other dialects. A conventional solution involves translating queries across dialects before training, using tools like SQLglot (Li et al., 2024b; Mao, 2023). However, this approach fails to leverage the unique capabilities of each dialect, as queries originally written for SQLite may not exploit the full potential of the target dialects. To overcome this, we propose a dialect-agnostic method for generating synthetic Text-to-SQL pairs for any database.
@@
SQL-GEN consists of a three-step pipeline. Initially, we begin with a small collection of seed SQL templates. These retain only the SQL keywords, abstracting away the specific database schema and values. Accompanying these templates, we provide dialect-specific tutorials that explain the usage of each SQL keyword across different dialects. In the first stage of our pipeline, we leverage a Large Language Model (LLM) to expand the seed templates, using the tutorials to adapt the keywords and function to various SQL dialects. In the second stage, these database-independent templates are populated with actual values and schema elements from any given database. The final stage involves a rigorous quality checks to ensure that the generated pairs accurately match each other.
@@
We perform comprehensive evaluations on the effectiveness of SQL-GEN. We assess the quality of synthetic samples by comparing them against both prior synthetic and human-annotated datasets. We construct synthetic datasets for three dialects and evaluate various LLMs (of sizes 7B-22B), trained on these pairs as well as other baselines. We show that all LLMs trained on our synthetic data exhibit a performance increase ranging from 4% to 27%, surpassing those trained on earlier synthetic and human-annotated data. In addition, for under-explored dialects of PostgreSQL and BigQuery, we did our evaluations on real-world data, specifically designed for PostgreSQL and BigQuery, models trained on our synthetic data consistently outperform others by a large margin, $\sim$7.5for BigQuery and $\sim$2.5for PostgreSQL. We also explore data augmentation for cross-domain Text-to-SQL scenario as another application of synthetic queries. By integrating synthetic queries with training samples from other databases, we enhance the model’s ability to adapt across domains. We test our proposed data augmentation method using the BIRD development set by combining synthetic and training data, aiming to improve performance on the target databases. We show consistent performance improvements of up to $\sim$5.7when fine-tuning on the augmented dataset.
@@
As noted earlier, each SQL dialect has distinct keywords and functions, rendering a model trained on a specific dialect uniquely specialized. However, a significant challenge arises when Text-to-SQL system users manage databases across multiple dialects, as deploying multiple dialect-specific models is computationally demanding. To overcome this, we introduced a novel method utilizing the Mixture of Experts (MoE) architecture (Fedus et al., 2022b; Riquelme et al., 2021; Jiang et al., 2024), compared this approach with existing techniques for merging models. Our method initializes the MLP layers of the MoE model using the MLP layers of the dialect-expert models, while the self-attention sub-layers are initialized using a two-by-two Spherical Linear Interpolation (SLERP) of self-attentions from the dialect experts. Additionally, to harness dialect-specific expertise effectively, we initialize the routers with hidden states corresponding to dialect-specific keywords. Our method demonstrated an improvement of approximately 2.5% in average performance over common model merging strategies and showed superior performance in SQLite and BigQuery dialects, outperforming the expert models by 0.68% and 7.25%, respectively. It also exceeded the performance of a MoE model initialized without our dialect-aware model merging, trained for the same number of steps.
@@ 2 Methodology
In this section, we first introduce the SQL-GEN pipeline, designed to generate high-quality, dialect-specific Text-to-SQL samples, as illustrated in Figure 2 and detailed in Algorithm 1. The generation of multi-dialect Text-to-SQL synthetic data addresses the critical issue of insufficient high-quality data for training models tailored to specific SQL dialects. However, since users of Text-to-SQL systems often work with databases across various dialects, serving models in a multi-dialect environment presents a unique challenge. Additionally, many SQL dialects adhere to standard SQL and share common syntaxes, making the concept of information sharing particularly compelling. To address these challenges, we propose a model merging method that combines dialect-specific models into a single, unified MoE model capable of serving multiple SQL dialects effectively.
@@ 2.1 Synthetic Text-to-SQL Data Generation
The initial step of SQL-GEN involves creating a pool of simple queries by extracting template question-SQL pairs. Building on this, we expand the templates using LLMs and dialect-specific tutorials rather than relying solely on extracted templates. After expanding these templates, each one is converted into an actual SQL query, and a corresponding question is generated by passing a sample database to the LLM. Subsequently, all generated question-SQL pairs, along with their execution results, undergo a quality-checking step to ensure they accurately match each other and effectively extract valuable information from the database. Throughout this process, we apply filtering to remove low-quality samples at each step.
@@
Similar to Wu et al. (2021); Yu et al. (2020), we extract SQL templates by abstracting all of the schema mentions in the queries from the Spider dataset to serve as a foundational pool for generating more diverse queries. Since the seed queries are initially written in SQLite, for the other dialects, we transpiled these queries to the other target dialects using the SQLGlot parser (Mao, 2023) before extracting their SQL query templates.
@@
The initial pool of query templates created in the previous step presents two main challenges. First, the extracted templates are derived from simple SQL queries, which are relatively basic compared to the queries found in other SQLite benchmarks like BIRD (Li et al., 2024b). Second, for dialects other than SQLite, the seed templates—originally designed for SQLite—do not utilize most of the dialect-specific SQL functions from other dialects. To address these, we expand the templates for each dialect using LLMs with in-context learning (Brown et al., 2020; Wei et al., 2022). To prepare the LLMs for template expansion, we first scrape online tutorials for each target dialect, focusing on the use of dialect-specific SQL functions and keywords. We then randomly select a seed template from the pool, pairing it with a random tutorial document about a dialect-specific keyword or function, and prompt the LLM to increase the complexity of the template, drawing inspiration from the document. To ensure the validity of the templates for all of the different dialects, we parse all generated SQL templates using SQLglot’s dialect-specific parser (Mao, 2023). Figure 9 demonstrates an examples of template expansion step for the BigQuery dialect. Additionally, Section A.10.1 provides the prompt that has been used for template expansion.
@@
After generating the SQL templates, our next step is to convert them into valid question-SQL pairs. For this process, we select a template along with a database schema with a random row from any given database. Random database rows are necessary since the LLM should be able to fill conditions with actual database values. We can source our database schemas from different datasets such as Spider or BIRD train/development databases. As these are originally in SQLite, we migrate these databases to each target dialect for dialects other than SQLite. We then pass the combination to an LLM, instructing it to integrate schema mentions into the templates and generate corresponding questions that align with the SQL queries. After generating the SQL queries, we apply heuristic-based semantic and syntactic filters to ensure the high quality of both the queries and questions. The specifics of these filters are detailed in the Section A.6. Additionally, Section A.10.2 includes the detailed prompt which is used in this step.
Data Quality Check:
To ensure high quality generation of question-SQL pairs, we present the question-SQL pairs alongside the first $K$ rows of their execution results over the database to an LLM. This LLM is tasked with verifying that the question and SQL pair match appropriately and that the question is free of ambiguity. To avoid repeating the same errors, we employ a different LLM, not used in previous steps, to act as the judge. Section A.4.1 provides a detailed analysis of the importance of utilizing a secondary LLM and highlights the importance of this step. Section A.10.3 provides the prompt that has been used for quality checking.
2.2 Dialect Experts Merging
With SQL-GEN, we can generate question-SQL pairs for various dialects and train corresponding dialect-specific models. However, in real-world scenarios, users often manage databases across different dialects, necessitating the deployment of multiple models, which can be resource-intensive. Additionally, while each dialect features unique keywords and functions, there is commonality across some SQL keywords across dialects. By merging these dialect experts into a single model, we facilitate sharing of common knowledge, addressing the challenge of serving multiple models simultaneously. We begin by detailing Spherical Linear Interpolation for model merging, and subsequently introduce our proposed method utilizing the Mixture of Experts (MoE) architecture.
SLERP:
Spherical Linear Interpolation, or SLERP, provides a method for achieving smooth transitions between two vectors. This technique ensures uniform motion and upholds the spherical space’s intrinsic geometric qualities. The process begins by normalizing the input vectors to have a magnitude of one. Subsequently, the angle between these vectors is determined. In instances where the vectors are almost col-linear, the method defaults to linear interpolation to optimize efficiency. For other cases, SLERP calculates scale factors that depend on the interpolation factor t and the angular separation of the vectors.
@@ 2.2 Dialect Experts Merging
MoE models are a sparse, decoder-only, transformer-based architecture, where each layer in this model contains more than one MLP blocks, or "experts," and a router network that selects experts to process each token at every position, combining their outputs. This architecture modifies the traditional MLP sub-layer within transformer blocks by replacing it with multiple experts, each having an independent set of parameters (Jiang et al., 2024; Fedus et al., 2022a).
In our proposed methodology, the MLP layers of each expert within the MoE model are initialized using the MLP sub-layers from models previously trained on distinct dialects. Concurrently, the shared self-attention layers incorporate merging strategies such as SLERP, as previously described. An important part of the MoE framework is its gating mechanism or routers. We propose to initialize these gates at each layer by averaging the hidden vectors of the dialect-specific keywords, derived from the training datasets of each model and based on the top $K$ most frequently occurring dialect-specific keywords in our generated question-SQL pairs. This process begins by cataloging all dialect-specific keywords from our generated SQL queries, sorting them by frequency, and selecting the top-k keywords. These keywords are then processed by the model, where the hidden representations from the self-attention sub-modules for all tokens of these keywords are used to initialize the gates. Although this method shows enhanced performance even without further training compared to some previous model merging techniques, we achieve better coordination among the sub-models by fine-tuning the MoE architecture on a mixed dataset from various dialects. Figure 3 illustrates one transformer block of our method for constructing an MoE model from three distinct dialect expert models.
@@
Figure 3: Our proposed method to initialize one transformer block of a MoE model from different dialect experts, exemplified here for Postgres, SQLite, and BigQuery dialects to create an all in one model to address all.
3 Experiment

### v2

@@ Abstract
Recent advances in Text-to-SQL have largely focused on the SQLite dialect, neglecting the diverse landscape of SQL dialects like BigQuery and PostgreSQL. This limitation is due to the diversity in SQL syntaxes and functions, along with the high cost of collecting and curating SQL-specific training data. To address this, we introduce SQL-GEN, a framework for generating high-quality synthetic training data for any SQL dialect, guided by readily available dialect-specific tutorials. SQL-GEN significantly improves cross-dialect Text-to-SQL performance, boosting execution accuracy by up to 20% over existing methods. This performance gain narrows the gap with models trained on large-scale human-annotated data. Furthermore, combining synthetic data from SQL-GEN with human-annotated data yields additional improvements of up to 5.6%. To unify multi-dialect capabilities within a single model, we propose a novel Mixture-of-Experts (MoE) initialization that leverages the shared knowledge across dialects. Our approach merges self-attention layers from dialect-specific models and initializes expert gates using dialect-specific keywords. This leads to a versatile model optimized for multiple SQL dialects, outperforming single-dialect models and significantly enhancing overall performance.
@@ 1 Introduction
The ability to seamlessly interact with databases using natural language is a long-sought goal in the field of human-computer interaction. Text-to-SQL systems, which translate natural language questions into executable SQL queries, are crucial for achieving this goal, bridging the gap between human communication and the structured language of databases (Androutsopoulos et al., 1995; Hristidis et al., 2003; Li & Jagadish, 2014). These systems have become essential components of modern conversational agents, empowering them to efficiently process complex queries within large-scale databases (Yu et al., 2019; Gu et al., 2022; Pérez-Mercado et al., 2023). Moreover, they serve as invaluable copilots for data science professionals, boosting productivity and enabling non-technical users to glean insights without SQL expertise (Li et al., 2023; Sun et al., 2023a, b; Wang et al., 2019).
SQL has been adopted by each database product (e.g. PostgreSQL, MySQL, and SQLite) to suit their specific needs. Despite their common foundations, these SQL dialects differ significantly in their syntax, functions, and capabilities, which even make the automated translation of queries across dialects a complex task that often requires human intervention (Zmigrod et al., 2024; Ngom & Kraska, 2024). Figure 1 exemplifies a question that can be answered with different SQL keywords across different dialects with their own unique keywords that are distinct from one another. Additionally in Section A.7, we provide some of the dialect specific keywords for BigQuery, PostgreSQL, and SQLite, which are not supported across all of them. In the realm of Text-to-SQL, most benchmarks are based on the SQLite dialect, chosen for its simplicity and self-contained nature (Li et al., 2024b; Yu et al., 2018b; Zhong et al., 2017; Chang et al., 2023; Gan et al., 2021). This dialect dependency poses a significant challenge, as models trained on SQLite-specific syntax are prone to generating erroneous queries in other dialects. A conventional solution involves translating queries across dialects before training, using tools like SQLglot parser or tools offered by cloud providers (Li et al., 2024b; Mao, 2023; Zmigrod et al., 2024). However, most of these tools are not 100% successful in translating the queries between dialects Zmigrod et al. (2024). For example, during the translation of the BIRD benchmark (Li et al., 2024b) from SQLite to BigQuery, approximately 20% of the queries encountered errors using the SQLGlot parser. Additionally, this approach fails to leverage the unique capabilities of each SQL dialect, as queries originally written for SQLite may not fully exploit the potential of the target dialects due to the absence of support for specific functions and keywords in the source dialect. For example, REGEX operations are supported in BigQuery but not in SQLite, so we cannot get this REGEX support by translating queries from SQLite.
@@
To address the aforementioned challenges, we propose SQL-GEN, a novel framework for dialect-agnostic synthetic generation of Text-to-SQL pairs for any database. SQL-GEN consists of a three major steps. Initially, we begin with a small collection of seed SQL templates. These retain only the SQL keywords, abstracting away the specific database schema and values. Accompanying them, we provide dialect-specific tutorials on each SQL keyword across different dialects. In the first stage, we leverage a Large Language Model (LLM) to expand the seed templates, using the tutorials to adapt the keywords and function to various SQL dialects. In the second stage, these database-independent templates are populated with actual values and schema elements from any given database. The final stage involves rigorous quality checks to ensure that the generated pairs accurately match each other. We conduct thorough evaluations to assess SQL-GEN’s effectiveness in teaching models new dialects, with a particular focus on dialect-specific features such as keywords. We assess the quality of synthetic samples by comparing them against both prior synthetic and human-annotated data. For three dialects, we evaluate various LLMs (of sizes 7B-22B), trained on these pairs as well as other baselines. We show that LLMs trained on SQL-GEN’s synthetic data exhibit a performance increase ranging from 4% to 27%, surpassing those trained on others. For under-explored dialects, we focus on evaluations on real-world data, specifically designed for them. We demonstrate that models trained SQL-GEN’s synthetic data consistently outperform others by a significant margin, approximately 7.5% on BigQuery and 2.5% on PostgreSQL dialect-specific datasets. This highlights the generalizability of SQL-GEN to unseen datasets due to its broad coverage. We also explore data augmentation for cross-domain Text-to-SQL as another use case of synthetic queries. By integrating synthetic queries with samples from other databases, we show improvements in models’ ability to adapt across domains. We show consistent performance improvements, up to $\sim$5.7%, fine-tuning with the augmented data.
@@
Each SQL dialect has distinct keywords and functions, rendering a model trained on a specific dialect uniquely specialized. A significant challenge arises when users manage databases across multiple dialects, as deploying multiple dialect-specific models can be computationally demanding. Moreover, we multi-dialect datasets might share common SQL characteristics, leading to some overlap in the features learned by the models. By merging models, we believe they can gain a deeper understanding of core features as they appear across multiple dialects. To overcome this, we introduce a novel method for utilizing the Mixture of Experts (MoE) architecture (Fedus et al., 2022b; Riquelme et al., 2021; Jiang et al., 2024). Specifically, our approach is based on initializing the MoE model using the layers of the dialect-expert models, while the sub-layers are initialized using a two-by-two Spherical Linear Interpolation (SLERP) of self-attentions from the dialect experts, as approaches for efficient merging. Additionally, to harness dialect-specific expertise effectively, we initialize the routers with hidden states corresponding to dialect-specific keywords. We demonstrate an improvement of 2.5% in average performance compared to other model merging approaches, as well as superior performance, outperforming the expert models up to 7.3%.
@@ 2 Methodology
We first introduce the SQL-GEN pipeline, designed to generate high-quality, dialect-specific Text-to-SQL samples, as illustrated in Figure 2 and detailed in Algorithm 1. The generation of multi-dialect Text-to-SQL synthetic data addresses the critical issue of insufficient high-quality data for training models tailored to specific SQL dialects. However, since users of Text-to-SQL systems often work with databases across various dialects, serving models in a multi-dialect environment presents a unique challenge. Additionally, many SQL dialects adhere to standard SQL and share common syntaxes, making the concept of information sharing particularly compelling. To this end, we propose a model merging method that combines dialect-specific models into a single, unified MoE model capable of serving multiple SQL dialects effectively.
@@ 2.1 Synthetic Text-to-SQL Data Generation
The initial step of SQL-GEN involves creating a pool of simple queries by extracting template question-SQL pairs. Building on this, we expand the templates using LLMs and dialect-specific tutorials, rather than relying solely on extracted templates. After expanding these templates, each one is converted into an actual SQL query, and a corresponding question is generated using an LLM with the sample database. Subsequently, all generated question-SQL pairs, along with their execution results, undergo a quality-checking step to ensure they accurately match each other and effectively extract valuable information from the database. Throughout this process, we apply filtering to remove low-quality samples at each step, to ensure the overall generated data would be high quality.
@@
Similar to (Wu et al., 2021; Yu et al., 2020), we extract SQL templates by abstracting all of the schema mentions in the queries from the Spider dataset to serve as a foundational pool for generating more diverse queries. As the seed queries are initially in SQLite, for the other dialects, we transpile these queries using the SQLGlot parser (Mao, 2023) before extracting their SQL query templates.
@@
The initial pool of query templates created in the previous step presents two main challenges. First, the extracted templates are derived from simple SQL queries, which are relatively basic compared to the queries found in other SQLite benchmarks like BIRD (Li et al., 2024b). Second, for dialects other than SQLite, the seed templates—originally designed for SQLite would not have complete coverage for all the dialect-specific SQL functions from other dialects. To address these, we expand the templates for each dialect using LLMs with in-context learning (Brown et al., 2020; Wei et al., 2022). To prepare the LLMs for template expansion, we first scrape online tutorials for each target dialect, focusing on the use of dialect-specific SQL functions and keywords. We then randomly select a seed template from the pool, pairing it with a random tutorial document about a dialect-specific keyword or function, and prompt the LLM to increase the complexity of the template, drawing inspirations from the document. To ensure the validity of the templates for all of the different dialects, we parse all generated SQL templates using dialect-specific parser (from SQLglot (Mao, 2023)). Figure 3 exemplifies this template expansion step for the BigQuery dialect. Additionally, Section A.9.1 provides the prompt that has been used for template expansion.
@@
After generating the SQL templates, our next step is to convert them into valid question-SQL pairs. For this process, we select a template along with a database schema with a random row from any given database. Random database rows are necessary since the LLM should be able to fill conditions with actual database values. The database schemas can be sourced from different datasets (e.g. publicly-available Spider or BIRD). As these are originally in SQLite, these databases are migrated to each target dialect for dialects other than SQLite. The combination is then passed to an LLM, instructing it to integrate schema mentions into the templates and generate corresponding questions that align with the SQL queries. After generating the SQL queries, heuristic-based semantic and syntactic filters are applied to ensure the high quality of both the queries and questions. The specifics of these filters are detailed in the Section A.6. Additionally, Section A.9.2 includes the detailed prompt which is used in this step.
Data Quality Checking:
To ensure high quality generation of question-SQL pairs, we present the pairs alongside the first $K$ rows of their execution results over the database to an LLM. This LLM is tasked with verifying that the pair match appropriately and that the question is free of ambiguity. To avoid repeating the same errors, we employ a different LLM, not used in previous steps, to act as the judge. Section A.4.1 provides a detailed analysis of the importance of utilizing a secondary LLM and highlights the importance of this step. Section A.9.3 provides the prompt that has been used for quality checking.
2.2 Dialect Experts Merging: Multi-dialects Mixture of Expert (MoE)
With SQL-GEN, we can generate question-SQL pairs for various dialects and train corresponding dialect-specific models. However, in real-world scenarios, users often manage databases across different dialects, necessitating the deployment of multiple models, which can come with practical challenges, including increased model serving costs and overhead of managing multiple checkpoints. Additionally, while each dialect features unique keywords and functions, there is commonality across some SQL keywords across dialects that can be exploited for cross-dialect information transfer. By merging these dialect experts into a single model, not only we mitigate the practical serving challenges, but we can improve the performance of each facilitating sharing of common knowledge. As model merging approaches, we introduce our proposed method utilizing the Mixture of Experts (MoE) architecture.
@@ 2.2 Dialect Experts Merging: Multi-dialects Mixture of Expert (MoE)
In a Mixture of Experts (MoE) model, each layer contains multiple MLP blocks, or “experts," and a router network selects specific experts to process each token at every position, combining their outputs. This architecture enhances the traditional MLP sub-layer within Transformer blocks by replacing it with multiple experts, each with its own set of parameters (Jiang et al., 2024; Fedus et al., 2022a). MoE-based LLMs route tokens to different experts, increasing modeling expressiveness without significantly increasing the compute budget, as only a subset of experts is activated for each token. With different Transformer-based expert models, we can combine them into a single MoE model that leverages the expert-specific MLP layers. By initializing the router to select the corresponding expert for each token, we can combine the knowledge of expert models by activating multiple experts and merging the self-attention layers. This approach aligns well with our setting, where we have dialect-specific expert models that already have prior knowledge of SQL syntax. For dialect-specific keywords, we can use the router to select the appropriate MLP layer, which allows us to integrate three models into a single one without the need to train a new model from scratch. Figure 4 illustrates one Transformer block with the proposed method for constructing an MoE model from three distinct dialect expert models.
@@
Figure 4: Our proposed method to initialize one Transformer block of a MoE model from different dialect experts, exemplified here for Postgres, SQLite, and BigQuery dialects to create an all in one model to address all. Objects in yellow demonstrate multi-dialect models
Multi-Dialect Routing:
An important aspect of the MoE framework is the routing (gating) mechanism. In MoE, the output for a given input $x$ is computed as a weighted sum of the expert networks’ outputs, with the weights determined by the gating network (Jiang et al., 2024). Given $n$ expert networks $\{E_{1},E_{i},...,E_{n}\}.$ The output is
$\sum_{i=0}^{n-1}G(x)_{i}E_{i}(x),$
where the gates’ outputs are determined based on the dot product of the input $x$ and the gate weights $W_{g}$ as follows:
$G(x)=\text{Softmax}(TopK(dot(x,W_{g}))).$
We propose to initialize these gates at each layer by averaging the hidden vectors of the dialect-specific keywords, derived from the training data of each model and based on the top $K$ most frequently occurring dialect-specific keywords from generated question-SQL pairs. This process begins by cataloging all dialect-specific keywords from our generated SQL queries, sorting them by frequency, and selecting the top-k keywords. These keywords are then processed by the model, where the hidden representations from the self-attention sub-modules for all tokens of these keywords are used to initialize the gates. The formula below provides the described method for initializing the gate weights:
$W_{g}(i)=\frac{1}{K}\sum_{k=1}^{K}\sum_{j=1}^{k_{t}}h_{k_{j}},$
where $W_{g}(i)$ is the ith column of the gate weight matrix corresponding to ith expert, $k_{t}$ is the number of tokens for the kth dialect keyword, and $h_{k_{j}}$ is the hidden representation of the j-th token of the k-th keyword. This approach increases the dot product between dialect-specific input keywords and their corresponding gate weight matrix columns, thereby boosting the weight for the dialect-specific expert. Although this method shows improved performance even without further training compared to other model merging approaches, we show superior joint modeling of the sub-models by further fine-tuning the MoE architecture on a mixed dataset from various dialects.
SLERP-based self-attention merging:
In our proposed approach, the MLP layers of each expert within the MoE model are initialized using the MLP sub-layers from models previously trained on distinct dialects. For the self-attention sub-layers of the MoE model, we employ Spherical Linear Interpolation (SLERP) (Goddard et al., 2024; Shoemake, 1985) to merge the initial weights of the self-attention layers (Key, Value, and Query projections) across multiple dialects. SLERP allows for smooth, non-linear transitions between two weight vectors while preserving the intrinsic geometric properties of the spherical space. The process begins by normalizing the weights of the Key, Value, and Query layers from different dialect models to unit magnitude, ensuring that they lie on the surface of a unit sphere. Once normalized, the angle ($\theta$) between the weight vectors is computed using the dot product. If the vectors are nearly col-linear (i.e., the dot product is close to 1), the merging process defaults to linear interpolation (LERP) for efficiency. Otherwise, SLERP estimates the scale factors based on the interpolation parameter $t$ and the angular separation between the vectors:
$\text{SLERP}(t,\mathbf{v}_{0},\mathbf{v}_{1})=\frac{\sin((1-t)\theta)}{\sin(\theta)}\mathbf{v}_{0}+\frac{\sin(t\theta)}{\sin(\theta)}\mathbf{v}_{1}.$
Here, $v_{0}$ and $v_{1}$ represent the normalized weight vectors of the models. By merging the self-attention weights through SLERP, we can smoothly integrate the knowledge from different dialect-specific models into the initialization of the MoE model’s self-attention layers, providing a more effective starting point for model training.
3 Experiments